OpenStudy (anonymous):

The first digit of a three digit number is one quarter of the second sight. The third digit is one quarter the value of the number formed by the first two digits. The sum of the three digits is 17. What is the number?

OpenStudy (fibonaccichick666):

so hint: can you write 4 equations to represent this paragraph?

OpenStudy (anonymous):

umm i got some but i dont know if they are correct, lets say the number = abc a=1/4b b=? c=1/4(a+b) a+b+c = 17

OpenStudy (fibonaccichick666):

ok lets only use 1 variable can you alter your equations to show me that?

OpenStudy (fibonaccichick666):

the equations are good though

OpenStudy (anonymous):

i dont know what u mean by altering the equations to 1 variable sorry

OpenStudy (fibonaccichick666):

(I just appreciate you trying) so instead of having a, b, c maybe only have b in one big eq

OpenStudy (anonymous):

oh okay! 1/4b + 1/4([1/4b]+b) + b = 17 like that?

OpenStudy (fibonaccichick666):

exactly can you solve for b now?

OpenStudy (anonymous):

i cant seem to figure it out :/

OpenStudy (fibonaccichick666):

ok let's do it together

OpenStudy (fibonaccichick666):

so I am going to start by distributing the 1/4 in the second term. can you write this step for me?

OpenStudy (anonymous):

do u mean 1/4b + 1/4([1/4b]+b)? sorry im a big confused

OpenStudy (fibonaccichick666):

ok give me a min and ill explain better

OpenStudy (anonymous):

okay thankyou so much

OpenStudy (fibonaccichick666):

http://www.educreations.com/lesson/view/for-oly-surf/7121943/?ref=link this should explain it

OpenStudy (fibonaccichick666):

did that help?

OpenStudy (anonymous):

thankyou yes it did, but when i figured out 25/16b=17, i got b=10.88 is that right?

OpenStudy (anonymous):

sorry i just saw the end of the video, what do u mean by plug it in for a and c?

OpenStudy (fibonaccichick666):

probably I didnt calculate it and by plug it in a mean just (1/4)b=a(first digit) so thats (1/4)(10.88)= bla bla bla

OpenStudy (anonymous):

i dont think its right, like my calculations, sorry i just cant get the 3 digit number dont worry if im being complicated i can ask my teacher

OpenStudy (fibonaccichick666):

ok h/o I'll redo the problem because you're right it doesn't make sense

OpenStudy (anonymous):

okay thankyou so much appreciate it

OpenStudy (fibonaccichick666):

ok I redid it and got the same answer are you sure you copied the problem correctly?

OpenStudy (anonymous):

what 3 digit number did u get?

OpenStudy (fibonaccichick666):

You can't get one, thats the problem lol

OpenStudy (anonymous):

yeah haha thought so! one sec il check if i wrote it correctly

OpenStudy (fibonaccichick666):

if you only took the first digit it would be 213..

OpenStudy (anonymous):

but it dosent equal 17 thats the problem

OpenStudy (anonymous):

The first digit of a three digit number is one quarter of the second digit. The third digit is one quarter the value of the number formed by the first two digits. The sum of the three digits is 17. What is the number? This was definitely the question

OpenStudy (fibonaccichick666):

ok well... I think they messed up lol let me try another approach

OpenStudy (anonymous):

dont worry if u cant find it thanks for the help anyway

OpenStudy (fibonaccichick666):

sorry I couldn't get it... I mean that is the answer. I guarantee it but that is so odd

OpenStudy (anonymous):

no worries thanks anyway

OpenStudy (fibonaccichick666):

@Hoa , can you double check this for me please?

OpenStudy (fibonaccichick666):

@nincompoop can you double check this for me?

OpenStudy (nincompoop):

xyz x/4=y z=x+y x+y+z = 17 sorry but I didn't use any mathematical methodologies to answer this… I came up with 287

OpenStudy (fibonaccichick666):

z=1/4(x+y)

OpenStudy (nincompoop):

not x+y it is 1/4 of xy The third digit is one quarter the value of the number formed by the first two digits. we are not using algebraic operations

OpenStudy (nincompoop):

1/4 of 28 is 7 our first two digits in 287 is 28

OpenStudy (fibonaccichick666):

OMG! formed as in put together... I'm such a nitwit usually I read formed as add

OpenStudy (nincompoop):

I thought I was going to embarrass myself in front of you…. I was sweating bullets (lol's)

OpenStudy (fibonaccichick666):

lol ha it's cool but i would have never thought of just making a number by putting them together. That works... although it's more of a riddle than a math problem in that case